Exercise 9.7 Accounting for Trade-ins (LO9-5)

Willis Bus Service traded in a used bus for a new one. The original cost of the old bus was $50,900. Accumulated depreciation at the time of the trade-in amounted to $35,400. The new bus cost $77,500 but Willis was given a trade-in allowance of $10,600.


a. What amount of cash did Willis have to pay to acquire the new bus?

b. Compute the gain or loss on the disposal for financial reporting purposes.
Answer #1

a.

Cost of new bus = $77,500

Trade in allowance = $10,600

Cash to be paid to acquire the new bus = Cost of new bus - Trade in allowance

= 77,500-10,600

= $66,900

b.

Original cost of old bus = $50,900

Accumulated depreciation = $35,400

Book value of old Bus = Original cost of old bus - Accumulated depreciation

= 50,900-35,400

= $15,500

Trade in allowance = $10,600

Loss on disposal of old bus = Book value of old Bus - Trade in allowance

= 15,500-10,600

= $4,900

Kindly comment if you need further assistance.

Thanks‼!